OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3.

OpenStudy (anonymous):

@ash2326 Any help?

OpenStudy (anonymous):

x1=-1 y1=6 m=-3 Now just plug into\[y-y_1=m(x-x_1)\]

OpenStudy (nubeer):

you have the slope = -3 and a point x1 = -1 ,y1 = 6 now apply this formula.. y-y1 =m(x-x1) just subsitute value you will get the equation

OpenStudy (anonymous):

I'm horrible with slope, so I kind of don't understand what you're trying to explain :( My options are: y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (anonymous):

@nubeer @Twis7ed ?

OpenStudy (anonymous):

@dpaInc ?

OpenStudy (nubeer):

ok first u need to remember the formula y-y1 = m(x-x1) in here m is the slope .. x1 is the x-coordinate and y1 is the y coordinate.. got this so far?

OpenStudy (anonymous):

Yes, kind of.

OpenStudy (nubeer):

ok now look which values u have.. u have m , x1 and y1, right? can u tell me what values does each of them have @unheard